OpenStudy (anonymous):

Determine the equation(s) of the vertical asymptote(s) of the function y=0.9^x/2x^2-3 plz help!

OpenStudy (anonymous):

set the denominator equal to zero, solve fo r\(x\)

OpenStudy (anonymous):

\[2x^2-3=0\\ 2x^2=3\\ x^2=\frac{3}{2}\\ x=\pm\sqrt{\frac{3}{2}}\] everything but that